The number of traffic accidents per week in a small city has a Poisson distribution with mean 

equal to 1.3. What is the probability of at least two accidents in 2 weeks?


1
Expert's answer
2022-05-22T23:33:09-0400

Let "X=" the number of traffic accidents per week: "X\sim Po(\lambda t)."

Given "\lambda=1.3, t=2."


"P(X\ge 2)=1-P(X=0)-P(X=1)"

"=1-\dfrac{e^{-2.6}(2.6)^0}{0!}-\dfrac{e^{-2.6}(2.6)^1}{1!}\approx0.73261512"